A rectangle is three times as long as it is wide.
The perimeter of the rectangle is 147 inches.
What is the length of the rectangle?
3X+x=147
4x/4 = 147/4
x = 36.75/ 2
= 18.37 x 3
=55.1
Did I do it correctly?

Answer by rothauserc(4718)

let x be the width, then 3x is the length
:
Perimeter of rectangle = 2*length + 2*width
:
2(3x) +2x = 147
:
6x +2x = 147
:
8x = 147
:
x = 18.375
:
************************************************
length of rectangle = 3 * 18.375 = 55.125 inches
:
check the answer
:
2 * 55.125 + 2 * 18.375 = 147
:
110.25 + 36.75 = 147
:
147 = 147
:
answer checks
************************************************